Find the directional derivative of the function f(x,y,z) = z4−x3y2 at P(1,-1,1) in the direction of the vector from P(1,-1,1) to the point Q(2,1,0). What is the maximum rate of change of f at the point P(1,-1,1) and in which direction
Answer #1

sod? Sind directonal Derivative 2 #(x,y,z) = 24-xy at point P(1,-1,1) Direction rector (1,1,1) to (2,1,0) ů = (2+) û + (1+1)sit It ou 4#(P) = -3(1%-1334-201214) ĝ xf = af ft af 94 at o 1,4,1) vp =-3x?y? - dzyjeg + 423% +42193 VA(P) = -3h + 2 9 +4Ã D